### v1

Abstract
Treewidth is an important structural graph parameter that quantifies how closely a graph resembles a tree-like structure. It has applications in many algorithmic and combinatorial problems. In this paper, we study treewidth of outer $k$-planar graphs – graphs admitting a convex drawing where all vertices lie on a circle and each edge crosses at most $k$ other edges. We also consider a more general class of outer min-$k$-planar graphs, which are graphs admitting a convex drawing where for every crossing of two edges at least one of these edges is crossed at most $k$ times.
Firman, Gutowski, Kryven, Okada and Wolff [GD 2024] proved that every outer $k$-planar graph has treewidth at most $21.5k+$ 2 and provided a lower bound of $k+2$ for even $k$. We establish a lower bound of $0.51.5k+$ 0.5 for every odd $k$. Additionally, they showed that every outer min-$k$-planar graph has treewidth at most $13k+$ 1. We improve this upper bound to $43\cdot{\left\lfloor 0.5k\right\rfloor}+$ 4.
Our approach also allows us to upper bound the separation number, a parameter closely related to treewidth, of outer min-$k$-planar graphs by $42\cdot{\left\lfloor 0.5k\right\rfloor}+$ 4. This improves the previous bound of $12k+$ 1 and achieves a bound with an optimal multiplicative constant.
keywords:
treewidth, outer $k$-planar graphs, outer min-$k$-planar graphs, separation number
1 Introduction
In this paper, we study classes of graphs admitting a convex drawing with bounded number of edge crossings. A convex drawing is a straight-line drawing with all vertices drawn on a common circle. Bannister and Eppstein [1, 2] proved that the treewidth of graphs admitting a convex drawing with at most $k$ crossings in total is bounded by a linear function of $\sqrt{k}$. For a fixed $k$, they also provided a linear time algorithm deciding if a given graph admits such drawing (using Courcelle’s theorem [6]). Another, studied before, class of graphs in this area is the class of outer $k$-planar graphs, that is graphs admitting a convex drawing, in which every edge crosses at most $k$ other edges. These graphs have bounded treewidth by a linear function of $k$, which was first proven by Wood and Telle [14, Proposition 8.5]. The authors of [5], also using Courcelle’s theorem, for a fixed $k$, showed a linear time algorithm testing whether given graph is maximal outer $k$-planar. Recently, Kobayashi, Okada and Wolff [11], for a fixed $k$, provided a polynomial time algorithm testing whether given graph is outer $k$-planar and proved that recognising outer $k$-planar graphs is XNLP-hard.
For disambiguation, we recall the definition of $k$-outerplanar graphs. A graph is outerplanar if it has a planar drawing with all vertices lying on the outer face. A graph is $1$-outerplanar when it is outerplanar. A graph is $k$-outerplanar for $k>1$ when it has a planar drawing such that after removing the vertices of the outer face, each of the remaining components is $(k-1)$-outerplanar.
We mainly study treewidth of outer $k$-planar graphs and outer min-$k$-planar graphs. A graph is outer min-$k$-planar if it admits a convex drawing, in which for every crossing of two edges at least one of these edges is crossed at most $k$ times. The authors of [8] proved that outer $k$-planar graphs have treewidth at most $21.5k+$ 2 and outer min-$k$-planar graphs have treewidth at most $13k+$ 1. To obtain these results, they showed that every outer $k$-planar graph admits a triangulation of the outer cycle such that every edge of the triangulation is crossed at most $k$ times by the edges of the graph. A similar property was proven for outer min-$k$-planar graphs.
Another property closely related to treewidth is the separation number of a graph. A separation of a graph $G$ is a pair $(A,B)$ of subsets of $V(G)$ such that $A\cup B=V(G)$ and there aren’t any edges between sets $A\setminus B$ and $B\setminus A$. The order of a separation is ${\left|A\cap B\right|}$. A separation is balanced if ${\left|A\setminus B\right|}\leq\frac{2}{3}{\left|V(G)\right|}$ and ${\left|B\setminus A\right|}\leq\frac{2}{3}{\left|V(G)\right|}$. The separation number of a graph $G$, denoted $\operatorname{sn}(G)$, is the minimum integer $a$ such that every subgraph of $G$ has a balanced separation of order at most $a$. Robertson and Seymour [12] proved that $\operatorname{sn}(G)\leq\operatorname{tw}(G)+1$ for every graph $G$. From the other side, Dvořák and Norin [7] showed that $\operatorname{tw}(G)\leq 15\operatorname{sn}(G)$. Recently, Houdrouge, Miraftab and Morin [9] provided a more constructive proof of an analogous inequality, but with a worse multiplicative constant.
Our contribution.
The authors of [8] proved that for every outer $k$-planar graph, its treewidth is at most $21.5k+$ 2. They also presented a lower bound of $k+2$ for every even $k$. We present an infinite family of outer $k$-planar graphs with treewidth at least $0.51.5k+$ 0.5, showing that the multiplicative constant $1.5$ in the upper bound cannot be improved; see Section 3.
We also improve the upper bounds for the treewidth and separation number of outer min-$k$-planar graphs. It was previously known that the treewidth of such graphs is at most $13k+$ 1 and the separation number is at most $12k+$ 1 [8]. We give an upper bound of $43\cdot{\left\lfloor 0.5k\right\rfloor}+$ 4 for treewidth; see Section 4, and an upper bound of $42\cdot{\left\lfloor 0.5k\right\rfloor}+$ 4 for the separation number; see Section 5. Both multiplicative constants are optimal, as the lower bounds for outer $k$-planar graphs also hold for outer min-$k$-planar graphs – namely, our lower bound of $0.51.5k+$ 0.5 for the treewidth and a lower bound of $k+2$ for the separation number presented in [8].
2 Preliminaries
Let $G$ be a graph. By $V(G)$ and $E(G)$ we denote the set of vertices and edges of $G$, respectively. For an edge, we use the compact notation $uv$, instead of $\left\{u,v\right\}$. For a directed edge, we stick to the standard notation $(u,v)$. Let $\deg(v)$ denote the degree of a vertex $v$ and $\Delta(G)$ be the maximum degree of a vertex of $G$.
For a graph $G$, a subgraph induced by a set $U\subseteq V(G)$, denoted $G[U]$, is a subgraph with vertex set $U$ and all edges of $G$ between the vertices of $U$. A spanning tree of a graph $G$ is a subgraph of $G$ containing all the vertices of $G$ that is a tree. By $\operatorname{dist}_{G}(v,w)$ we denote the distance (i.e. the length of the shortest path) between $v$ and $w$ in a graph $G$. For any tree $T$ rooted at vertex $r$, we define the depth of a vertex $v$ as $0pt_{T}(v)=\operatorname{dist}_{T}(r,v)$. We may omit subscripts if they are clear from the context.
A tree decomposition $\mathcal{T}=(T,B)$ of a graph $G$ is a collection of nodes, called bags, $\{B_{x}:x\in V(T)\}$ indexed by the vertices of a tree $T$. Every bag is a subset of $V(G)$ with following properties satisfied:
  1. 1.
    for every vertex $v\in V(G)$ the set $\{x:v\in B_{x}\}$ induces a non-empty subtree in $T$;
  2. 2.
    for every edge $uv\in E(G)$ there exists a bag containing both $u$ and $v$.
The width of a given tree decomposition is the size of the largest bag minus one. The treewidth of a graph $G$, denoted by $\operatorname{tw}(G)$, is the minimum width of any tree decomposition of $G$.
A set $\mathcal{B}$ of non-empty subsets of $V(G)$ is a bramble if:
  1. 1.
    for every $X\in\mathcal{B}$ the induced subgraph $G[X]$ is connected;
  2. 2.
    for every $X_{1},X_{2}\in\mathcal{B}$ the induced subgraph $G[X_{1}\cup X_{2}]$ is connected. That means, each $X_{1},X_{2}\in\mathcal{B}$ share a common vertex or there exists an edge of $G$ incident to both $X_{1}$ and $X_{2}$.
A hitting set of a bramble is a set of vertices with non-empty intersection with every element of $\mathcal{B}$. The order of a bramble is the size of its smallest hitting set. The bramble number of a graph $G$, denoted by $\operatorname{bn}(G)$, is the maximum order of any bramble of $G$.
The following result by Seymour and Thomas shows the relation between bramble number and treewidth.
Theorem 2.1 (Seymour and Thomas, [13]).
For every graph $G$, $\operatorname{tw}(G)=\operatorname{bn}(G)-1$.
We say that a graph $G$ is a minor of a graph $H$, if $G$ can be obtained from $H$ by a sequence of vertex deletions, edge deletions or edge contractions. Edge contraction of an edge $uv$ is an operation that replaces vertices $u$ and $v$ with a new vertex adjacent to every vertex other than $u$ and $v$ that was adjacent to $u$ or $v$. It is a well-known fact that if $G$ is a minor of $H$ then $\operatorname{tw}(G)\leq\operatorname{tw}(H)$. The proof of this fact can be found in [3].
In the remainder of this section, we introduce some notation and simple observations regarding drawings. A convex drawing of a graph $G$ is a straight-line drawing where the vertices of $G$ are placed on different points of a circle in the cyclic order $(v_{1},\ldots,v_{n})$. We say that an edge $v_{i}v_{j}$ with $i<j$ crosses an edge $v_{i^{\prime}}v_{j^{\prime}}$ with $i^{\prime}<j^{\prime}$ if either $1\leq i<i^{\prime}<j<j^{\prime}\leq n$ or $1\leq i^{\prime}<i<j^{\prime}<j\leq n$. We consider only such convex drawings that no three edges pass through the same point. An outer $k$-planar drawing of a graph is a convex drawing such that every edge crosses at most $k$ other edges. An outer min-$k$-planar drawing of a graph is a convex drawing such that for every crossing of two edges at least one of these edges crosses at most $k$ other edges.
An outer $k$-planar graph $G$ is maximal outer $k$-planar if for every $e\in V^{2}(G)\setminus E(G)$ the graph $G+e$ is not outer $k$-planar.
{observation}
Let $G$ be a maximal outer $k$-planar graph with at least three vertices. Then, in every outer $k$-planar drawing of $G$, the outer face is bounded by a simple cycle.
Proof 2.2.
Consider an outer $k$-planar drawing $\Gamma$ of $G$. Let $u$ and $v$ be consecutive vertices in the cyclic order defined by $\Gamma$. Suppose that $uv\not\in E(G)$. Notice that the graph $G+uv$ has an outer $k$-planar drawing defined by the same cyclic order as $\Gamma$, which contradicts the maximality of $G$.
A graph $G$ is expanded outer $k$-planar if $G$ is an outer $k$-planar graph with $\Delta(G)\leq 3$ and its outer face is bounded by a simple cycle in some outer $k$-planar drawing of $G$.
{observation}
Every outer $k$-planar graph $G$ is a minor of an expanded outer $k$-planar graph $G^{\prime}$.
Proof 2.3.
Let us assume that $G$ is maximal outer $k$-planar. Now, in order to obtain $G^{\prime}$ from $G$, we perform the following transformation to every vertex $v$ of $G$ with $\deg(v)\geq 4$. The transformation is depicted in Figure 1. Let $w_{0},w_{1},\ldots,w_{s},w_{s+1}$ be all neighbors of $v$ in clockwise order, with edges $vw_{0}$ and $vw_{s+1}$ adjacent to the outer face of $G$. We replace $v$ with a path $v_{1},\ldots,v_{s}$, put it on the outer face of $G$ in counter clockwise order, in the place of $v$. We connect this path to vertices $w_{0}$ and $w_{s+1}$ by adding edges $v_{1}w_{0}$ and $v_{s}w_{s+1}$. Finally, for every $1\leq i\leq s$, we add an edge $v_{i}w_{i}$ that corresponds to an edge $vw_{i}$ in the original graph. It is easy to see that $G$ is a minor of $G^{\prime}$ and the ordering of corresponding edges in $G^{\prime}$ matches the one in $G$. The crossings in the resulting graph naturally correspond to the crossings in the original graph.
Figure 1: The transformation described in Section 2.
The vertices $v_{1},\ldots,v_{s}$ defined in the proof are called images of $v$ and $v$ is the origin of these vertices, denoted $\operatorname{org}(v_{i})=v$. If the transformation was not performed for some vertex $v$ of $G$, i.e. $\deg(v)\leq 3$, then $v$ is an image and origin of itself.
We remark that the analogous definitions and Observations 2, 2 hold for outer min-$k$-planar graphs. Since adding edges increases neither the treewidth nor the separation number, we are interested in the properties of maximal graphs. Also, taking a minor does not increase the treewidth, so we work with expanded graphs when bounding treewidth.
3 Lower bound for treewidth of outer $k$-planar graphs
In this section, we construct an infinite family of outer $k$-planar graphs with treewidth at least $0.51.5k+$ 0.5. This improves the previous lower bound of $k+2$ that was presented in [8]. We begin by defining the necessary graphs.
For positive integers $m$ and $n$, let $X_{m,n}$ denote the grid of $m$ rows and $n$ columns, i.e. a graph with
$V(X_{m,n})=\left\{x_{i,j}:1\leq i\leq m,1\leq j\leq n\right\}\text{ and }E(X_{m,n})=\left\{x_{i,j}x_{k,l}:{\left|i-k\right|}+{\left|j-l\right|}=1\right\}.$
For a positive integer $k$, consider $Q_{k}$ – a copy of grid $X_{2k,2k}$ and $R_{k}$ – a copy of $X_{2k(k+1),k}$. Let $v_{i,j}$, for $1\leq i,j\leq 2k$, be a vertex in $i$-th row and $j$-th column of $Q_{k}$, and $u_{i,j}$, for $1\leq i\leq 2k(k+1)$, $1\leq j\leq k$, be a vertex in $i$-th row and $j$-th column of $R_{k}$. Let $G_{k}$ be a graph such that $V(G_{k})=V(Q_{k})\cup V(R_{k})$ and
$E(G_{k})=E(Q_{k})\cup E(R_{k})\cup\left\{v_{i,2k}u_{(i-1)(k+1)+j,1}:1\leq i\leq 2k,1\leq j\leq k+1\right\};$
see Figure 2. For $1\leq i\leq 2k(k+1)$, let $i$-th extended row of $G_{k}$ be the union of $i$-th row of $R_{k}$ and $\lceil\frac{i}{k+1}\rceil$-th row of $Q_{k}$. Notice that each row of $Q_{k}$ is contained in $k+1$ extended rows and the graph induced by each extended row is a path.
The graph $G_{k}$ was previously defined by Kammer and Tholey in [10] as an example of tightness of the upper bound for the treewidth of $k$-outerplanar graphs. They used the cops and robber game to establish the lower bound for treewidth of $G_{k}$. Below, we present a proof using brambles.
Figure 2: The graph $G_{k}$, for $k=3$, with a subgraph of $\mathcal{B}_{1}$ colored green and a subgraph of $\mathcal{B}_{2}$ colored orange.
Theorem 3.1 (Kammer and Tholey, [10]).
For every $k\geq 1$, $\operatorname{tw}(G_{k})=3k-1$.
Proof 3.2.
Notice that the drawing of $G_{k}$ in Figure 2 is $k$-outerplanar. By the fact that $k$-outerplanar graphs have treewidth at most $13k-$ 1 (Bodlaender, [4]), we get $\operatorname{tw}(G_{k})\leq 3k-1$.
To prove that $\operatorname{tw}(G_{k})\geq 3k-1$, we will construct a bramble of order $3k$. Then, using Theorem 2.1, we will get $\operatorname{tw}(G_{k})\geq 3k-1$. Let $\mathcal{B}_{1}$ be a family consisting of every subgraph of $G_{k}$ that is a union of an extended row of $G_{k}$ and a column of $Q_{k}$. Let $\mathcal{B}_{2}$ be a family consisting of every subgraph of $G_{k}$ that is a union of a row of $R_{k}$ and a column of $R_{k}$. The set $\mathcal{B}=\mathcal{B}_{1}\cup\mathcal{B}_{2}$ forms a bramble of $G_{k}$, as each subgraph in $\mathcal{B}$ is connected and every two subgraphs have at least one common vertex.
Consider any hitting set $S$ of $\mathcal{B}$. Let $q$ and $r$ be the number of vertices of $S$ in $V(Q_{k})$ and in $V(R_{k})$, respectively. Now, we would like to show that $|S|=q+r\geq 3k$. Notice that $r\geq k$, as otherwise there is a row and a column of $R_{k}$ not containing any element of $S$, and thus there is an element of $\mathcal{B}_{2}$ not hit by $S$.
If $q\geq 2k$ then $q+r\geq 3k$. Otherwise, let $q=2k-l$ for some positive integer $l$. Now, we can find at least $l$ columns and at least $l$ rows of $Q_{k}$ not intersecting $S$. These $l$ rows are contained in $l(k+1)$ extended rows. Every one of them has to intersect $S$ at some vertex of $R_{k}$, because otherwise we can find a column of $Q_{k}$ and an extended row not intersecting $S$ that form an element of $\mathcal{B}_{1}$. The extended rows restricted to $R_{k}$ are pairwise disjoint so we have $r\geq l(k+1)$. Summing up, we get $q+r\geq 2k-l+l(k+1)=2k+lk\geq 2k+k=3k$, which concludes the proof.
Let $F_{k}$ be the following modification of $G_{k}$ depicted in Figure 3. We set $\ell(i)=(k-i)(k+1)$ for $1\leq i\leq k$, and $\ell(i)=(i-k-1)(k+1)$ for $k+1\leq i\leq 2k$. For every $1\leq i\leq 2k$, we remove every edge between $v_{i,2k}$ and any vertex of the grid $R_{k}$. We add a path $Z_{i}$ of length $\ell(i)$ on vertices $z_{i,0}z_{i,1}\ldots{}z_{i,\ell(i)}$. We add a path $W_{i}$ of length $k$ on vertices $w_{i,1}w_{i,2}\ldots{}w_{i,k+1}$. We connect $v_{i,2k}$ with $Z_{i}$ by adding an edge $v_{i}z_{i,0}$; $Z_{i}$ with $W_{i}$ by adding an edge $z_{i,\ell(i)}w_{i,k+1}$ for $1\leq i\leq k$ or an edge $z_{i,\ell(i)}w_{i,1}$ for $k+1\leq i\leq 2k$. Finally, we connect $W_{i}$ with $R_{k}$ by adding an edge $w_{i,j}u_{(i-1)(k+1)+j,1}$ for every $11\leq j\leq k+$ 1.
To see that $G_{k}$ is a minor of $F_{k}$ it is enough to contract, for every $1\leq i\leq 2k$, vertex $v_{i,2k}$ with vertices of paths $Z_{i}$ and $W_{i}$. Since taking a minor does not increase the treewidth, we obtain the following corollary.
Figure 3: The graph $F_{k}$ – a modification of the graph $G_{k}$, for $k=3$.
Corollary 3.3.
For every $k\geq 1$, $\operatorname{tw}(F_{k})\geq 3k-1$.
Theorem 3.4.
The graph $F_{k}$ has an outer $(2k-1)$-planar drawing for every $k\geq 1$.
Proof 3.5.
We describe an outer $(2k-1)$-planar drawing of $F_{k}$ as depicted in Figure 4. We call the set of vertices $\{v_{i,j}:1\leq i\leq k,1\leq j\leq 2k\}$ the upper part of $Q_{k}$. The other vertices of $Q_{k}$ are called the lower part of $Q_{k}$. We define a cyclic order of the vertices of $F_{k}$ by arranging them in a clockwise direction from some selected starting point on a circle.
(a) Overview of the outer $(2k-1)$-planar drawing of $F_{k}$.
(b) Drawing of the upper part of the grid $Q_{k}$.
(c) Drawing of the groups $k,k-1,\ldots,1$ connected to the upper part of $Q_{k}$.
(d) Drawing of the grid $R_{k}$.
Figure 4: Fragments of the outer $(2k-1)$-planar drawing of $F_{k}$, where $k=3$.
First, we put vertices from the upper part of $Q_{k}$, in the column-by-column order (see Figure 4(b)):
$v_{k,1},\ldots,v_{2,1},v_{1,1},$
$v_{k,2},\ldots,v_{2,2},v_{1,2},$
$\vdots$
$v_{k,2k},\ldots,v_{2,2k},v_{1,2k}.$
After that, we put vertices $z_{k,0},\ldots,z_{2,0},z_{1,0}$, in this order.
We divide the remaining vertices of paths $Z_{i}$ and $W_{i}$, for $1\leq i\leq k$, into $k$ groups, as follows. The $i$-th group contains vertices of $W_{i}$, and if $i\geq 2$, it also includes vertices $z_{a,b}$ for $1\leq a<i$ and $(k-i)(k+1)<b\leq(k-i+1)(k+1)$. Next, on the drawing, we respectively put the groups of indices $k,k-1,\ldots,1$. We arrange the vertices in the $i$-th group, for $2\leq i\leq k$, in the order (see Figure 4(c)):
$w_{i,k+1},z_{i-1,(k-i)(k+1)+1},z_{i-2,(k-i)(k+1)+1},\ldots,z_{1,(k-i)(k+1)+1},$
$w_{i,k},z_{i-1,(k-i)(k+1)+2},z_{i-2,(k-i)(k+1)+2},\ldots,z_{1,(k-i)(k+1)+2},$
$\vdots$
$w_{i,1},z_{i-1,(k-i+1)(k+1)},z_{i-2,(k-i+1)(k+1)},\ldots,z_{1,(k-i+1)(k+1)}.$
The group of index 1 has vertices arranged in the order: $w_{1,k+1},w_{1,k},\ldots,w_{1,1}$.
Next, we put the vertices of $R_{k}$ in the row-by-row order (see Figure 4(d)):
$u_{1,1},u_{1,2},\ldots,u_{1,k},$
$u_{2,1},u_{2,2},\ldots,u_{2,k},$
$\vdots$
$u_{2k(k+1),1},u_{2k(k+1),2},\ldots,u_{2k(k+1),k}.$
The vertices of $F_{k}$ that are not placed yet are in the lower part of $Q_{k}$ or in paths $Z_{i}$, $W_{i}$, for $k+1\leq i\leq 2k$. We arrange them in a counter clockwise direction from the starting point and place them between the starting point and the vertices of $R_{k}$. The order is symmetric, with respect to the starting point, to the one used to arrange the upper part of $Q_{k}$ and paths $Z_{i}$, $W_{i}$ for $1\leq i\leq k$. Every vertex $v_{i,j}$, where $k+1\leq i\leq 2k$ and $1\leq j\leq 2k$, is placed symmetrically to $v_{2k-i+1,j}$. Vertices $z_{i,j}$, where $k+1\leq i\leq 2k$ and $0\leq j\leq\ell(i)$, are placed symmetrically to $z_{2k-i+1,j}$, and vertices $w_{i,j}$, where $k+1\leq i\leq 2k$ and $11\leq j\leq k+$ 1, symmetrically to $w_{2k-i+1,k-j+2}$. The symmetrical drawing of the $i$-th group, for every $1\leq i\leq k$, forms the group of index $12k-i+$ 1.
Now, we will show that every edge crosses at most $12k-$ 1 other edges, partitioning them into several types.
  1. 1.
    The “column” edges in the upper or lower part of $Q_{k}$ – edges $v_{i,j}v_{i+1,j}$, for $1\leq i\leq 2k-1,i\neq k$ and $1\leq j\leq 2k$. These edges cross no other edges.
  2. 2.
    The “column” edges between the upper and the lower part of $Q_{k}$ – edges $v_{k,j}v_{k+1,j}$, for $1\leq j\leq 2k$. Each of these edges crosses $k-1$ edges of type 3 from the upper part of $Q_{k}$, and $k-1$ edges from the lower part. The edge $v_{k,1}v_{k+1,1}$ crosses no other edges.
  3. 3.
    The “row” edges of $Q_{k}$ – edges $v_{i,j}v_{i,j+1}$, for $1\leq i\leq 2k$ and $11\leq j\leq 2k-$ 1. Each of these edges crosses $2(k-1)$ edges of types 3, 4 and additionally at most one edge of type 2.
  4. 4.
    Each edge $v_{i,2k}z_{i,0}$, for $1\leq i\leq 2k$, crosses $2(k-1)$ edges either of type 3 or edges incident to vertices $z_{j,0}$.
  5. 5.
    Each edge $z_{i,\ell(i)}w_{i,k+1}$, for $1\leq i\leq k$, crosses exactly $2(i-1)$ edges incident to vertices $z_{i-1,0},\ldots,z_{1,0}$. Symmetrically, each edge $z_{i,\ell(i)}w_{i,1}$, for $k+1\leq i\leq 2k$, also crosses at most $2(k-1)$ edges.
  6. 6.
    Each edge $z_{i,0}z_{i,1}$, for $1\leq i\leq 2k$, $i\not\in\{k,k+1\}$, crosses exactly $2(k-2)$ other edges from the paths $Z_{j}$ or incident to vertices $z_{j,0}$; and exactly 3 edges incident to $w_{k,k+1}$ or $w_{k+1,1}$.
  7. 7.
    Every other edge from paths $Z_{i}$ crosses at most $2(k-2)$ edges from other paths $Z_{j}$ and at most 3 edges incident to some vertex $w_{a,b}$.
  8. 8.
    Each edge from paths $W_{i}$ crosses at most $2(k-1)$ edges from paths $Z_{j}$.
  9. 9.
    Each edge $w_{i,j}u_{(i-1)(k+1)+j,1}$, for $1\leq i\leq 2k$ and $11\leq j\leq k+$ 1, crosses at most $k-1$ edges from paths $Z_{a}$ and at most $k-1$ edges from $R_{k}$.
  10. 10.
    The “row” edges of $R_{k}$ – edges $u_{i,j}u_{i,j+1}$, for $1\leq 2k(k+1)$ and $11\leq j\leq k-$ 1. They cross no other edges.
  11. 11.
    The “column” edges of $R_{k}$ – edges $u_{i,j}u_{i+1,j}$, for $11\leq 2k(k+1)-$ 1 and $1\leq j\leq k$. Each of these edges crosses at most $2(k-1)$ other edges of this type and at most one edge of type 9.
Theorem 3.6.
For every odd positive integer $k$, there exists an outer $k$-planar graph $G$ with $\operatorname{tw}(G)\geq 1.5k+0.5$.
Proof 3.7.
By Theorem 3.4, the graph $F_{\frac{k+1}{2}}$ is outer $k$-planar and by Corollary 3.3 has treewidth at least $0.53\frac{k+1}{2}-1=1.5k+$ 0.5.
4 Upper bound for treewidth of outer min-$k$-planar graphs
In this section, we upper bound the treewidth of outer min-$k$-planar graphs. We improve the previous bound of $13k+$ 1 presented in [8] to $43\cdot{\left\lfloor 0.5k\right\rfloor}+$ 4. We begin by introducing required notation.
For an outer min-$k$-planar graph $G$ with a given drawing $\Gamma$, we define a crossing graph $G_{C}$ as a graph, in which the vertex set is a union of $V(G)$ and all crossing points of the edges of $G$. We say that a vertex $w\in V(G_{C})$ lies on an edge $uv\in E(G)$ if $w$ is an endpoint of $uv$ or the crossing point corresponding to $w$ belongs to the segment that is a drawing of the edge $uv$ in $\Gamma$. Graph $G_{C}$ contains an edge between two vertices if and only if they are consecutive vertices lying on the drawing of some edge of $G$. Observe that $G_{C}$ is a planar graph. We say that an edge $xy\in E(G_{C})$ lies on an edge $uv\in E(G)$ if both $x$ and $y$ lie on $uv$ in $\Gamma$. Furthermore, we say that a vertex $v\in V(G_{C})$ is outer if it is adjacent to the outer face of $G_{C}$. Otherwise, $v$ is an inner vertex. As we consider only maximal graphs $G$, the outer vertices of $G_{C}$ are exactly the vertices of $G$. By $f_{o}$ we will denote the outer face of $G_{C}$.
For a planar graph $G$, denote $G^{*}$ as the graph dual to $G$. By $f^{*}\in V(G^{*})$ we denote the vertex dual to the face $f$ of $G$ and by $e^{*}\in E(G^{*})$ we denote the edge dual to an edge $e\in E(G)$. We remark that $G^{*}$ can be drawn on the drawing of $G$ in a way that $f^{*}$ is on the face $f$ and the drawing of $e^{*}$ is a curve that passes through the edge $e$ and faces corresponding to the endpoints of $e^{*}$.
The following lemma shows a bijection between a spanning tree $T$ of a planar graph $G$ and a spanning tree of $G^{*}$, that we denote by $T^{*}=\operatorname{dual}(T)$. We also use notation $\operatorname{dual}(T^{*})$ for $T$.
Lemma 4.1 (Folklore).
Let $T$ be a spanning tree of a planar graph $G$. Then $T^{*}$ with $V(T^{*})=V(G^{*})$ and $E(T^{*})=\left\{e^{*}:e\in E(G)\setminus E(T)\right\}$ is a spanning tree of $G^{*}$.
The next lemma proves that there exists a spanning tree preserving shortest paths from a given vertex. Such tree can be found via a BFS algorithm.
Lemma 4.2 (Folklore).
Let $G$ be a graph and let $r$ be a vertex of $G$. Then there exists a spanning tree $T$ of $G$ rooted at $r$ such that $0pt_{T}(v)=\operatorname{dist}_{G}(r,v)$ for every vertex $v$ of $G$.
Lemma 4.3.
Let $G$ be an expanded outer min-$k$-planar graph with its crossing graph $G_{C}$. Then $\operatorname{dist}(f^{*},f_{o}^{*})\leq{\left\lfloor 0.5k\right\rfloor}+1$ for every vertex $f^{*}\in V(G_{C}^{*})$.
Proof 4.4.
Let $f$ be a non-outer face of $G_{C}$. If $f$ is adjacent to $f_{o}$ then $\operatorname{dist}(f^{*},f^{*}_{o})=1$. Otherwise, let $v$ be a vertex of $G_{C}$ adjacent to $f$. As $G$ is expanded, the vertex $v$ is inner, so it lies on an edge $e$ of $G$ that crosses at most $k$ other edges. Let $v_{0},v_{1},\ldots,v_{s},v_{s+1},\ldots,v_{s+t+1}$ be all vertices lying on $e$, listed in the consecutive order, where $v_{s}=v$ and $v_{s+1}$ is a neighbor of $v$ that is adjacent to $f$. We may assume that $s\leq t$, i.e. $v_{s}$ is closer to an endpoint of the edge $e$ than $v_{s+1}$ to the other endpoint of $e$. Notice that at most $k+2$ vertices lie on $e$ (two endpoints and at most $k$ crossing points), so $s+t+2\leq k+2$. Together with the previous inequality, this implies $s\leq 0.5k$. The number $s$ is an integer so $s\leq{\left\lfloor 0.5k\right\rfloor}$.
We inductively define a sequence $w_{s},w_{s-1},\ldots,w_{0}$ of vertices. Vertex $w_{s}$ is the neighbor of $v_{s}$ that is adjacent to $f$ and not lying on $e$. For every $i=s-1,\ldots,0$, the vertex $w_{i}$ is one of the two neighbors of $v_{i}$ not lying on $e$ such that $w_{i},v_{i},v_{i-1}$ and $w_{i-1}$ are adjacent to the same face of $G_{C}$. Let $e_{i}$ denote the edge $v_{i}w_{i}$. Now, notice that the path formed by the edges $e_{s}^{*},\ldots,e_{0}^{*}$ connects $f^{*}$ with $f_{o}^{*}$ ($e_{0}$ is adjacent to $f_{o}$). So, $\operatorname{dist}(f^{*},f_{o}^{*})\leq s+1\leq{\left\lfloor 0.5k\right\rfloor}+1$.
Let $v$ be an inner vertex of $G_{C}$ with neighbors $w_{1},w_{2},w_{3},w_{4}$ in clockwise order. As we forbid common crossing points of three edges of $G$, every such vertex has degree 4. We subdivide the vertex $v$ by replacing it with two vertices $v_{1}$ and $v_{2}$ connected with an edge, and adding edges $v_{1}w_{1}$, $v_{1}w_{2}$ and $v_{2}w_{3}$, $v_{2}w_{4}$. We fix a planar embedding of the new graph by drawing $v_{1}$, $v_{2}$ very close to where $v$ was drawn. We say that vertices $v_{1}$ and $v_{2}$ lie on the same edges as the vertex $v$. Moreover, we say that the edge $v_{1}v_{2}$ in an auxiliary edge. After subdividing a vertex, every edge of $G_{C}$ has an edge corresponding to it and every face of $G_{C}$ corresponds to a new one in a natural way. Also, the dual graph has one additional edge dual to $v_{1}v_{2}$.
By $G_{S}$ we denote the subdivided crossing graph – the graph $G_{C}$ with all inner vertices subdivided. See Figure 5 for an example. Observe that there is a one-to-one correspondence between $V(G_{C}^{*})$ and $V(G_{S}^{*})$. Further, every edge of $G_{C}^{*}$ has a corresponding edge of $G_{S}^{*}$. The following lemma shows how we can preserve the properties of a spanning tree $T_{C}$ of $G_{C}$ and a spanning tree $\operatorname{dual}(T_{C})$ of the dual graph after subdividing vertices of $G_{C}$.
Lemma 4.5.
Let $G$ be an expanded outer min-$k$-planar graph with its subdivided crossing graph $G_{S}$. Then there exists a spanning tree $T_{S}$ of $G_{S}$ and a spanning tree $T_{S}^{*}=\operatorname{dual}(T_{S})$ of $G_{S}^{*}$ rooted at $f_{o}^{*}$, such that $10pt(f^{*})\leq{\left\lfloor 0.5k\right\rfloor}+$ 1 for every vertex $f^{*}\in V(G_{S}^{*})$ and $E(T_{S})$ contains all auxiliary edges of $G_{S}$.
Proof 4.6.
Let $G_{C}$ be a crossing graph of $G$. By Lemmas 4.2 and 4.3 there exists a spanning tree $T_{C}^{*}$ of $G_{C}^{*}$ that is rooted at $f_{o}^{*}$ and with maximal depth of a vertex being at most ${\left\lfloor 0.5k\right\rfloor}+1$. Let $T_{C}=\operatorname{dual}(T_{C}^{*})$ be spanning tree of $G_{C}$.
Let $G_{S}$ denote the graph obtained from $G_{C}$ by subdividing every inner vertex. After this transformation, let $T_{S}^{*}$ be a tree constructed of edges corresponding to the edges of $T_{C}^{*}$. Clearly, $T_{S}^{*}$ is a spanning tree of the graph $G_{S}^{*}$. The spanning tree $T_{S}=\operatorname{dual}(T_{S}^{*})$ of $G_{S}$ contains all auxiliary edges of $G_{S}$, because none of the duals of auxiliary edges are in $E(T_{S}^{*})$, as they were not in $E(T_{C}^{*})$.
Now, we are ready to prove the main result of this section.
Theorem 4.7.
Let $G$ be an outer min-$k$-planar graph. Then $\operatorname{tw}(G)\leq 3\cdot{\left\lfloor 0.5k\right\rfloor}+4$.
Proof 4.8.
By Section 2 we may assume that $G$ is an expanded outer min-$k$-planar graph. Let $G_{S}$ be the subdivided crossing graph of $G$. By Lemma 4.5 there exists a spanning tree $T_{S}$ of $G_{S}$ and a spanning tree $T_{S}^{*}=\operatorname{dual}(T_{S})$ of $G_{S}^{*}$ rooted at $f_{o}^{*}$, such that $10pt(f^{*})\leq{\left\lfloor 0.5k\right\rfloor}+$ 1 for every vertex $f^{*}\in V(G_{S}^{*})$ and $E(T_{S})$ contains all auxiliary edges of $G_{S}$.
(a) The graph $G_{C}$ with its spanning tree $T_{C}$ colored black and a spanning tree $T_{C}^{*}$ of $G_{C}^{*}$ in red.
(b) The graph $G_{S}$ with its spanning tree $T_{S}$ colored black and a spanning tree $T_{S}^{*}$ of $G_{S}^{*}$ in red.
Figure 5: Drawings of an example graphs with their spanning trees and spanning trees of the dual graphs, the vertex $f_{o}^{*}$ is missing on both figures.
Lets orient every edge of $G$. Edges adjacent to the outer face are oriented clockwise, the other edges are oriented arbitrarily. Observe that every vertex of $G$ has at most two incoming edges.
Now, we construct the tree decomposition $\mathcal{T}=(T_{S},B)$ of the graph $G$. The bags of $\mathcal{T}$ are indexed by the vertices of $T_{S}$. We put vertices of $G$ into bags using the following rules.
  1. 1.
    For every outer vertex $v\in V(G_{S})$, we put $v$ in the bag $B_{v}$.
  2. 2.
    For every oriented edge $(x,y)$ of $G$, we put $x$ in the bag $B_{y}$.
  3. 3.
    For every inner vertex $v\in V(G_{S})$, lying on the edge $(x,y)$ of $G$, we put $x$ in the bag $B_{v}$.
  4. 4.
    For every vertex $v\in V(G_{S})$, every non-outer face $f$ adjacent to $v$ and every edge $e^{*}$ belonging to the path from $f^{*}$ to $f_{o}^{*}$ in $T_{S}^{*}$ that is dual to the edge $e$ of $G_{S}$ lying on the edge $(x,y)$ of $G$, we put $x$ in the bag $B_{v}$.
Notice that, in rule 4, the edge $e$ is not in $E(T_{S})$, so it is not an auxiliary edge, which implies that it is lying only on a single edge of $G$.
For every edge $(x,y)$ of $G$, by rules 1 and 2, the bag $B_{y}$ contains both $x$ and $y$. Also, every vertex of $G$ is present in some bag. So to prove that $\mathcal{T}$ is a proper tree decomposition of $G$ it is enough to prove that for every vertex $x\in V(G)$ the set $\{w:x\in B_{w}\}$ induces a connected subtree in $T_{S}$.
Lets fix a vertex $x$ and an edge $(x,y)$ of $G$. Let $e=uv$ be an edge lying on $(x,y)$ such that $e\not\in E(T_{S})$. We assume that $e^{*}=f_{1}^{*}f_{2}^{*}$ and $10pt_{T_{S}^{*}}(f_{2}^{*})=0pt_{T_{S}^{*}}(f_{1}^{*})+$ 1. Let $T_{e}^{*}$ be a subtree of $T_{S}^{*}$ induced on all descendants of $f_{2}^{*}$, including $f_{2}^{*}$. Define $F_{e}=\left\{f:f^{*}\in V(T_{e}^{*})\right\}$ and let $\operatorname{boundary}(F_{e})$ be the set of vertices adjacent to some face in $F_{e}$. Notice that, by rule 4, $x$ is put into all bags indexed by $\operatorname{boundary}(F_{e})$. The set $\operatorname{boundary}(F_{e})$ induces a connected subgraph of $T_{S}$, i.e. $T_{S}[\operatorname{boundary}(F_{e})]$ is connected, containing both $u$ and $v$. Observe that the bags of $\mathcal{T}$, into which we put $x$ by rule 4, are exactly the bags of vertices of $T_{S}$ that are in $\operatorname{boundary}(F_{e})$ for some edge $e\not\in E(T_{S})$ lying on some edge $(x,y)$ of $G$.
By rules 1, 2 and 3, $x$ is contained in all bags $B_{w}$ such that $w$ lies on $(x,y)$ for some edge $(x,y)$ of $G$. We claim that the vertices indexing these bags, together with the vertices indexing bags we put $x$ into by rule 4, form a connected subgraph of $T_{S}$. To see that, we show that for every vertex $w$ such that $x\in B_{w}$, $w$ is connected to $x$ by a walk in $T_{S}$ such that bags indexed by the vertices of this walk contain $x$.
If $w$ lies on an edge $(x,y)$ of $G$ then, in order to construct this walk, we start at vertex $w$. We iterate over consecutive edges lying on $(x,y)$ between $w$ and $x$, starting at the edge incident to $w$. If given edge $e$ is in $E(T_{S})$ than we extend the walk by $e$. Otherwise, $e\not\in E(T_{S})$. As $T_{S}[\operatorname{boundary}(F_{e})]$ is connected and every bag of a vertex in $\operatorname{boundary}(F_{e})$ contains $x$, we can extend the walk by some path in $T_{S}[\operatorname{boundary}(F_{e})]$ connecting the endpoints of $e$.
If $w$ is in the set $\operatorname{boundary}(F_{e})$ for some $e$ lying on $(x,y)$, than we begin the walk with a path contained in $\operatorname{boundary}(F_{e})$ between $w$ and an endpoint $v$ of $e$. We extend this walk by a walk between $v$ and $x$, which existence we have already proven.
Next, we bound the size of bags in $\mathcal{T}$. Consider an inner vertex $v$ of $T_{S}$. It lies on exactly two edges of $G$, so by rule 3 we put two vertices into $B_{v}$. Also, $v$ is adjacent to three non-outer faces of $G_{S}$. For every such face $f$ and every edge $e^{*}$ belonging to the path from $f^{*}$ to $f_{o}^{*}$ in $T_{S}^{*}$, by rule 4, we put one vertex into $B_{v}$. Every such path has at most ${\left\lfloor 0.5k\right\rfloor}+1$ edges. So ${\left|B_{v}\right|}\leq 2+3\cdot\left({\left\lfloor 0.5k\right\rfloor}+1\right)$. Now, let $v$ be an outer vertex of $T_{S}$. By rules 1 and 2, the bag $B_{v}$ contains $v$ and at most two other endpoints of the edges incoming to $v$ in $G$. Also, $v$ is adjacent to two non-outer faces of $G_{S}$. Thus, we derive a bound ${\left|B_{v}\right|}\leq 3+2\cdot\left({\left\lfloor 0.5k\right\rfloor}+1\right)$. The width of the constructed tree decomposition is at most
$\max\left(2+3\cdot\left({\left\lfloor 0.5k\right\rfloor}+1\right),3+2\cdot\left({\left\lfloor 0.5k\right\rfloor}+1\right)\right)-1=2+3\cdot\left({\left\lfloor 0.5k\right\rfloor}+1\right)-1=3\cdot{\left\lfloor 0.5k\right\rfloor}+4.$
5 Separation number of outer min-$k$-planar graphs
The inequality $\operatorname{sn}(G)\leq\operatorname{tw}(G)+1$ bounding the separation number holds for every graph $G$. We remark that Theorem 4.7 directly implies $\operatorname{sn}(G)\leq 3\cdot{\left\lfloor 0.5k\right\rfloor}+5$ for every outer min-$k$-planar graph $G$. By carefully choosing some bag $B_{x}$ of the tree decomposition we can construct such balanced separation $(C,D)$ satisfying $C\cap D=B_{x}$. To establish a better upper bound, first we prove a general lemma showing how from a tree decomposition satisfying some additional properties, we can obtain a balanced separation $(C,D)$ such that $C\cap D=B_{x}\cap B_{y}$ for some two neighboring vertices $x,y$ of the tree decomposition.
Lemma 5.1.
Let $\mathcal{T}=(T,B)$ be a tree decomposition of a graph $G$. Assume that $\Delta(T)\leq 3$ and every vertex $v\in V(G)$ is in at least two bags of $\mathcal{T}$. Let $a$ be an integer such that ${\left|B_{x}\cap B_{y}\right|}\leq a$ for any edge $xy\in E(T)$. Then $G$ has a balanced separation of order at most $a$.
Proof 5.2.
For every edge $xy\in E(T)$, after removing it from $T$ we obtain two connected components $C_{x}$ and $C_{y}$ of $T$ such that $x\in V(C_{x})$ and $y\in V(C_{y})$. We define $S_{x,y}=\bigcup_{v\in V(C_{x})}B_{v}$ and $S_{y,x}=\bigcup_{v\in V(C_{y})}B_{v}$. It is a well known fact that the pair $(S_{x,y},S_{y,x})$ is a separation of $G$ of order ${\left|S_{x,y}\cap S_{y,x}\right|}={\left|B_{x}\cap B_{y}\right|}\leq a$.
We claim that there exists an edge $xy\in E(T)$ such that $(S_{x,y},S_{y,x})$ is a balanced separation of $G$. Suppose the contrary. Now, we orient every edge of $T$. For every $xy\in E(T)$ it holds that ${\left|S_{x,y}\setminus S_{y,x}\right|}>\frac{2}{3}n$ or ${\left|S_{y,x}\setminus S_{x,y}\right|}>\frac{2}{3}n$, where $n={\left|V(G)\right|}$. If the first inequality holds then we orient $xy$ as $(y,x)$, in the other case as $(x,y)$. Also, notice that ${\left|S_{x,y}\setminus S_{y,x}\right|}>\frac{2}{3}n$ is equivalent to ${\left|S_{y,x}\right|}<\frac{1}{3}n$.
The tree $T$ with oriented edges in an acyclic graph, so there exists a vertex in $T$ such that all edges incident to $x$ are oriented towards $x$. Let $y_{1},\ldots,y_{d}$, where $d\leq 3$, be all neighbors of $x$ in $T$. We have ${\left|S_{y_{i},x}\right|}<\frac{1}{3}n$. Also $\bigcup_{1\leq i\leq d}S_{y_{i},x}=\bigcup_{v\in V(T)\setminus\{x\}}B_{v}=V(G)$, as every vertex of $G$ is in at least two bags of $\mathcal{T}$. We obtain the following inequalities
${\left|V(G)\right|}={\left|\bigcup_{1\leq i\leq d}S_{y_{i},x}\right|}\leq\sum_{1\leq i\leq d}{\left|S_{y_{i},x}\right|}<d\cdot\frac{1}{3}n\leq n,$
which gives a contradiction.
Now, we are ready to upper bound the separation number of outer min-$k$-planar graphs.
Theorem 5.3.
Let $G$ be an outer min-$k$-planar graph. Then $\operatorname{sn}(G)\leq 2\cdot{\left\lfloor 0.5k\right\rfloor}+4$.
Proof 5.4.
The class of outer min-$k$-planar graphs is closed under taking subgraphs. So it is enough to find a balanced separation of order at most $42\cdot{\left\lfloor 0.5k\right\rfloor}+$ 4 for every maximal outer min-$k$-planar graph $G$. Let $H$ be an expanded outer min-$k$-planar graph obtained from $G$ by Section 2. By Theorem 4.7, there exists a tree decomposition $\mathcal{T}^{\prime}=(T_{S},B^{\prime})$ of $H$, where $T_{S}$ is a spanning tree of subdivided graph $H$. From the proof of Theorem 4.7, it follows that $\Delta(T_{S})\leq 3$ and every vertex $v\in V(H)$ is in at least two bags of $\mathcal{T}^{\prime}$ (there is an oriented edge $(v,w)$ in $H$, so $v\in B^{\prime}_{v}$ and $v\in B^{\prime}_{w}$).
We construct a tree decomposition $\mathcal{T}=(T_{S},B)$ of $G$ with $B_{x}=\left\{\operatorname{org}(v):v\in B_{x}^{\prime}\right\}$. Every vertex $v\in V(G)$ is in at least two bags of $\mathcal{T}$ as every image of $v$ is in at least two bags of $\mathcal{T}^{\prime}$. Every edge $vw\in E(G)$ is realised in some bag of $\mathcal{T}$, because in $H$ there is an edge corresponding to $vw$ between an image of $v$ and an image of $w$. To prove that for every vertex $v$ of $G$ the bags of $\mathcal{T}$ containing $v$ are spanning a connected subtree of $T_{S}$, lets denote all images of $v$ by $v_{1},\ldots,v_{s}$, in the consecutive order. As $H$ is maximal, for every $1\leq i<s$ there is an edge $v_{i}v_{i+1}$ in $E(H)$. So the two subtrees of $T_{S}$ induced by bags of $\mathcal{T}^{\prime}$ containing $v_{i}$, and bags of $\mathcal{T}^{\prime}$ containing $v_{i+1}$ share a common vertex. Bags containing $v$ in $\mathcal{T}$ are spanning a connected subtree of $T_{S}$, because this subtree is a union of subtrees spanned by the images of $v$. So $\mathcal{T}$ is a proper tree decomposition of $G$.
We say that a vertex $v$ was put into a bag $B_{x}^{\prime}$ of $\mathcal{T}^{\prime}$ due to rule 4 of constructing the tree decomposition being applicable to a vertex $v$ and a face $f$ adjacent to $x$, if there exists an edge $e$ lying on an edge $(v,w)$ of $H$ such that $e^{*}$ belongs to the path in $T_{S}^{*}$ between $f_{o}^{*}$ and $f^{*}$. Now, we want to show that for every edge $xy\in E(T_{S})$ we have ${\left|B_{x}\cap B_{y}\right|}\leq 2\cdot{\left\lfloor 0.5k\right\rfloor}+4$. Let $f_{1}$ and $f_{2}$ be the faces of $H_{S}$ adjacent to $xy$.
{claim*}
If $v\in B_{x}\cap B_{y}$ then there is an image $v_{t}$ of $v$ such that $xy$ lies on an edge $(v_{t},w)$ of $H$ or $v_{t}$ was put into both $B_{x}^{\prime}$ and $B_{y}^{\prime}$ due to rule 4 of constructing $\mathcal{T}^{\prime}$ being applicable to $v_{t}$ and face $f_{1}$ or $f_{2}$.
{claimproof}
If $\deg(x)=3$ in $H_{S}$, let $f_{x}$ be the face of $H_{S}$, different from $f_{1}$ and $f_{2}$ such that $f_{x}$ is adjacent to the vertex $x$. Similarly, if $\deg(y)=3$ in $H_{S}$, let $f_{y}$ be the face of $H_{S}$, different from $f_{1}$ and $f_{2}$ such that $f_{y}$ is adjacent to the vertex $y$. Assume that $v\in B_{x}\cap B_{y}$, but no images of $v$ were put into $B_{x}^{\prime}$ and $B_{y}^{\prime}$ due to the reasons in the claim statement. So there exist $v_{i_{x}}$ and $v_{i_{y}}$ that are, not necessarily distinct, images of $v$ such that $v_{i_{x}}\in B_{x}^{\prime}$ and $v_{i_{y}}\in B_{y}^{\prime}$. For $t\in\left\{x,y\right\}$, vertex $v_{i_{t}}$ was put in $B_{t}^{\prime}$ because
  1. 1.
    $t$ lies on an edge $(v_{i_{t}},w_{t})$ of $H$ such that $xy$ is not lying on $v_{i_{t}}w_{t}$; or
  2. 2.
    when $\deg(t)=3$ in $H_{S}$, rule 4 of constructing $\mathcal{T}^{\prime}$ is applicable to vertex $v_{i_{t}}$ and face $f_{t}$, i.e. there exists an edge $(v_{i_{t}},w_{t})$ of $H$ and an edge $e_{t}$ of $H_{S}$ lying on $v_{i_{t}}w_{t}$ such that $e_{t}^{*}$ belongs to the path between $f_{t}^{*}$ and $f_{o}^{*}$ in $T_{S}^{*}$.
Now, we draw a curve $\mathcal{C}$ on the drawing of $H_{S}$. $\mathcal{C}$ consists of the drawing of the edge $xy$ and the drawing of an arc of the outer face between $v_{i_{x}}$ and $v_{i_{y}}$ that contains only images of $v$ (images of $v$ are spanning a single arc of the outer face). Next, we add to $\mathcal{C}$ curves connecting $v_{i_{x}}$ with $x$ and $v_{i_{y}}$ with $y$. For $t\in\left\{x,y\right\}$, to determine how to draw these curves, we case over the reason $v_{i_{t}}$ was put in $B_{t}^{\prime}$, in the order as listed above.
  1. 1.
    We draw along the drawing of $(v_{i_{t}},w_{t})$, starting at vertex $v_{i_{t}}$ and ending at vertex $t$.
  2. 2.
    Let $p_{t}$ be a path between $f_{t}^{*}$ and $f_{o}^{*}$ in $T_{S}^{*}$. We draw along the drawing of $(v_{i_{t}},w_{t})$, staring at vertex $v_{i_{t}}$ and ending at the crossing point with the drawing of $p_{t}$. We continue along $p_{t}$ till vertex $f_{t}^{*}$. Finally, we connect vertices $f_{t}^{*}$ and $t$ with a segment.
Figure 6: Drawing of an example curve $\mathcal{C}$.
Note that if $x=v_{i_{x}}$ and $y=v_{i_{y}}$, then $\mathcal{C}$ is degenerated to the arc between $v_{i_{x}}$ and $v_{i_{y}}$, implying that $xy$ connects two consecutive images of $v$ – contradiction. Otherwise, we claim that one of the closed regions induced by $\mathcal{C}$ contains $f_{1}$ or $f_{2}$. Indeed, $\mathcal{C}$ follows edges of $H_{S}$ and edges of $T_{S}^{*}$, but cannot contain $f_{1}^{*}$ nor $f_{2}^{*}$, because then rule 4 of constructing $\mathcal{T}^{\prime}$ would be applicable to vertex $v_{i_{x}}$ or $v_{i_{y}}$ and face $f_{1}$ or $f_{2}$. The segments between $f_{t}^{*}$ and $t$ does not intersect $f_{1}$ nor $f_{2}$. We may assume that $f_{1}$ is contained inside a closed region induced by $\mathcal{C}$. Consider a path $p_{1}$ between $f_{1}^{*}$ and $f_{o}^{*}$ in $T_{S}^{*}$. As $f_{1}^{*}$ is inside $\mathcal{C}$ and $f_{o}^{*}$ is outside $\mathcal{C}$, drawing of $p_{1}$ has to intersect $\mathcal{C}$. We consider where the first intersection point is located.
  • Path $p_{1}$ cannot intersect $e$ nor the segments between $f_{t}^{*}$ and $t$.
  • If $p_{1}$ intersects an edge $(v_{i_{t}},w_{t})$ then rule 4 of constructing $\mathcal{T}^{\prime}$ is applicable to $v_{i_{t}}$ and $f_{1}$.
  • If $p_{1}$ intersects $p_{t}$ then $p_{1}$ follows along $p_{t}$ up to the intersection point with $(v_{i_{t}},w_{t})$, so the previous case applies.
  • If $p_{1}$ intersects the arc of the outer face between $v_{i_{x}}$ and $v_{i_{y}}$ then it has to intersect an edge $(v_{r},v_{r+1})$, where $v_{r}$ and $v_{r+1}$ are consecutive images of $v$ on the outer face. So rule 4 of constructing $\mathcal{T}^{\prime}$ is applicable to $v_{r}$ and $f_{1}$.
In each case we obtained a contradiction.
We proved that if $v\in B_{x}\cap B_{y}$ then there is an image $v_{t}$ of $v$ such that $xy$ lies on an edge $(v_{t},w)$ of $H$ or $v_{t}$ was put into both $B_{x}^{\prime}$ and $B_{y}^{\prime}$ due to rule 4 of constructing the tree decomposition $\mathcal{T}^{\prime}$ being applicable to vertex $v_{t}$ and face $f_{1}$ or $f_{2}$. Notice that $xy$ lies on at most two edges of $H$ (two if $xy$ is an auxiliary edge, one otherwise) and each of the paths from $f_{o}^{*}$ to $f_{1}^{*}$ or $f_{2}^{*}$ in $T_{S}^{*}$ has at most ${\left\lfloor 0.5k\right\rfloor}+1$ edges. So ${\left|B_{x}\cap B_{y}\right|}\leq 2+2\cdot({\left\lfloor 0.5k\right\rfloor}+1)=2\cdot{\left\lfloor 0.5k\right\rfloor}+4$. By Lemma 5.1 applied to $\mathcal{T}$ we get that $G$ has a balanced separation of order at most $42\cdot{\left\lfloor 0.5k\right\rfloor}+$ 4.
To give a lower bound we define a graph called stacked prism. A stacked prism $Y_{m,n}$ is an $m\times n$ grid with additional edges connecting the vertices of the first and the last row that are in the same column. The $Y_{m,n}$ has an outer $(2n-2)$-planar drawing, thus also an outer min-$(2n-2)$-planar drawing. In the cyclic order of the drawing, we place rows consecutively, one after another. The edges from rows cross no other edges and the edges from columns cross exactly $22n-$ 2 other edges. Authors of [8] showed that for every number $n$ and sufficiently large even number $m$, $\operatorname{sn}(Y_{m,n})=2n$. This leads to the following theorem.
Theorem 5.5.
For every even number $k$, there exists an outer min-$k$-planar graph $G$ such that $\operatorname{sn}(G)=k+2$.
We remark that the multiplicative constant of 1 in the upper bound given in Theorem 5.3 is tight, as it matches that of the lower bound in Theorem 5.5.